# Quadrilateral Particle Arrangement within Shocks in a Two-Dimensional Dusty Plasma

### v1

@@ I Introduction
The microscopic properties of shocks in 2D dusty plasma have been studied using the method of Voronoi diagrams, with an input of particle-position data obtained from MD simulations. In Ref. [12], the authors studied melting, which was reported to occur when the moving exciter had a speed exceeding a threshold, as determined by Voronoi diagrams that were examined qualitatively and were also used to quantify defect fraction. A leap of the defect ratio, which can indicate the occurrence of a phase transition from the solid to liquid states, was observed. In Ref. [13], a postshock region of 2D crystal, with both liquid and solid coexisting regions, was observed when the exciter speed was sufficiently high. For this phenomenon, again Voronoi diagrams were used to examine disorder and quantify defect fraction.
@@ II.1 Experiment
Compressional shocks were generated by continuously moving an exciter at a controlled velocity. The exciter was a nickel wire, shown in the Supplemental Material  [^1] The Supplemental Material for this preprint will be published, along with the article itself, in a future journal publication.. This exciter wire was aligned perpendicular to its direction of motion but parallel to the 2D microparticle cloud. Runs were performed at five different speeds of the exciter at various speeds, ranging from 50.8 to 101.6 mm/s all above the sound speed of 19 mm/s. In dimensionless units, this range was from 2.5 to 5.0 $a\omega_{pd}$, significantly exceeding the exciter speeds in the simulation of Sun et al. [15].
@@ II.2 Polylgon analysis
The Voronoi method has been used to study 2D melting with experimental data since at least the 1987 colloid experiments of Ref. [37]. The Voronoi algorithm is to form a polygon having its edges being the perpendicular bisectors of the bonds. In an equilibrium state of a 2D crystal, these Voronoi polygons will be hexagons, which is to say the local geometry has six-fold coordination. Five- and seven-sided polygons, and any other polygon that is not six-fold, are identified as a disclination, which is a defect. This kind of defect analysis has an essentially binary outcome: the vertex at the center of a polygon is either six-fold or it is not. Sufficient displacements of particles from the ground-state arrangement can result in a pair of polygons changing from hexagons to a five-seven pair of Voronoi cells. Sometimes only a tiny displacement of a single particle can result in such a change, from non-defect to defect. This binary classification, defect vs non-defect, is a disadvantage of the Voronoi method.
@@ IV Discussion
An analysis method that had not previously been applied to 2D shocks allowed us to find that within a shock, a hexagonal microstructure rearranged into a quadrilateral microstructure that had not be reported before. This analysis method was the Glaser-Clark polygon analysis, which we compared to the more common Voronoi analysis method. The data that we analyzed were from runs of the 2D dusty plasma shock experiment of Ref. [9]. An advantage of the Glaser-Clark polygons that we found was that it easily reveals quadrilateral structures, while Voronoi polygons are not well suited for this purpose, as they are mostly 5, 6 or 7-sided. Another advantage of Glaser-Clark polygons, which was reported previously by Ruhunusiri and Goree [27], is the gradation of defects, classified by the number of sides of the polygon increases as the disorder in the microstructure becomes greater.

### v2

@@ I Introduction
The microscopic properties of shocks in 2D dusty plasma have been studied using the method of Voronoi diagrams, with an input of particle-position data obtained from MD simulations. Sun and Feng [12] studied melting, which was reported to occur when the moving exciter had a speed exceeding a threshold, as determined by Voronoi diagrams that were examined qualitatively and were also used to quantify defect fraction. A leap of the defect ratio, which can indicate the occurrence of a phase transition from the solid to liquid states, was observed. Qiu et al. [13] studied a postshock region of 2D crystal, with both liquid and solid coexisting phases, which was observed when the exciter speed was sufficiently high. For this phenomenon, again Voronoi diagrams were used to examine disorder and quantify defect fraction.
@@ II.1 Experiment
Compressional shocks were generated by continuously moving an exciter at a controlled velocity. The exciter was a nickel wire, shown in the Supplemental Material  [^1] See Supplemental Material at [URL will be inserted by publisher]. This exciter wire was aligned perpendicular to its direction of motion but parallel to the 2D microparticle cloud. Runs were performed at five different speeds of the exciter at various speeds, ranging from 50.8 to 101.6 mm/s all above the sound speed of 19 mm/s. In dimensionless units, this range was from 2.5 to 5.0 $a\omega_{pd}$, significantly exceeding the exciter speeds in the simulation of Sun et al. [15].
@@ II.2 Polylgon analysis
The Voronoi method has been used to study 2D melting with experimental data since at least the 1987 colloid experiments of Ref. [37]. The Voronoi algorithm is to form a polygon having its edges being the perpendicular bisectors of the bonds. In an equilibrium state of a 2D crystal, these Voronoi polygons will be hexagons, which is to say the local geometry has six-fold coordination. Five- and seven-sided polygons, and any other polygon that is not six-fold, are identified as a disclination, which is a defect. This kind of defect analysis has an essentially binary outcome: the vertex at the center of a polygon is either six-fold or it is not. Sufficient displacements of particles from the ground-state arrangement can result in a pair of polygons changing from hexagons to a five-seven pair of Voronoi cells. Sometimes only a tiny displacement of a single particle can result in such a change, from non-defect to defect. This binary classification, defect vs non-defect, is a limitation of the Voronoi method.
@@ IV Discussion
An analysis method that had not previously been applied to 2D shocks allowed us to find that within a shock, a hexagonal microstructure rearranged into a quadrilateral microstructure that had not be reported before. This analysis method was the Glaser-Clark polygon analysis, which we compared to the more common Voronoi analysis method. The data that we analyzed were from runs of the 2D dusty plasma shock experiment of Ref. [9]. An advantage of the Glaser-Clark polygons that we found was that it easily reveals quadrilateral structures, while Voronoi polygons are not well suited for this purpose, as they are mostly 5, 6 or 7-sided. Another advantage of Glaser-Clark polygon anlaysis is that a gradually increasing disorder in the microstructure is quantified by a gradual increase in the number of sides of the polygons [27].